# Identify As A Human Does: A Pathfinder of Next-Generation Anti-Cheat Framework for First-Person Shooter Games

### v1

@@ 6.4 Ablation Study Results
Table 6: Hawk detection result with different Task-Specified Threshold Optimizer settings.
CT1 Optimizer2 Validation Test
TP TN FP FN Accuracy Recall AUC-ROC NPV OEI3 TP TN FP FN Accuracy Recall AUC-ROC NPV OEI3
Aimbot F1-Score 448 7,305 392 416 0.906 0.519 0.734 0.946 5.000 548 6,928 1235 336 0.826 0.620 0.734 0.954 3.000
Accuracy 269 7,528 169 595 0.911 0.311 0.645 0.927 5.640 305 7,888 275 579 0.906 0.345 0.656 0.932 5.014
AUC-ROC 648 4,603 3,094 216 0.613 0.750 0.674 0.955 1.639 691 4,364 3,799 193 0.559 0.782 0.658 0.958 1.508
$\text{Acc}_{\textit{rc}>0.75}$ 652 5,321 2,376 212 0.698 0.755 0.723 0.962 2.052 672 5,365 2,798 212 0.667 0.760 0.709 0.962 1.907
$\text{Acc}_{\textit{rc}>0.8}$ 700 4,611 3,086 164 0.620 0.810 0.705 0.966 1.769 716 4,700 3,463 168 0.599 0.810 0.693 0.965 1.693
$\text{Acc}_{\textit{rc}>0.85}$ 736 3,790 3,907 128 0.529 0.852 0.672 0.967 1.519 761 3,875 4,288 123 0.512 0.861 0.668 0.969 1.495
$\text{Acc}_{\textit{rc}>0.9}$ 778 2,859 4,838 86 0.425 0.900 0.636 0.971 1.333 787 2,955 5,208 97 0.414 0.890 0.626 0.968 1.301
Wallhack F1-Score 576 8,313 432 391 0.915 0.596 0.773 0.955 5.481 651 8,040 1,209 374 0.846 0.635 0.752 0.956 3.352
Accuracy 374 8,582 163 593 0.922 0.387 0.684 0.935 6.543 167 9,158 91 858 0.908 0.163 0.577 0.914 5.932
$\text{Acc}_{\textit{rc}>0.7}$ 677 7,898 847 290 0.883 0.700 0.802 0.965 4.304 803 7,506 1,743 222 0.809 0.783 0.797 0.971 3.071
$\text{Acc}_{\textit{rc}>0.75}$ 727 7,589 1,156 240 0.856 0.752 0.810 0.969 3.759 859 6,964 2,285 166 0.761 0.838 0.795 0.977 2.675
$\text{Acc}_{\textit{rc}>0.8}$ 776 7,186 1,559 191 0.820 0.802 0.812 0.974 3.251 583 8,280 969 442 0.863 0.569 0.732 0.949 3.574
$\text{Acc}_{\textit{rc}>0.85}$ 822 6,712 2,033 145 0.776 0.850 0.809 0.979 2.831 871 6,810 2,439 154 0.748 0.850 0.793 0.978 2.579
$\text{Acc}_{\textit{rc}>0.9}$ 871 5,755 2,990 96 0.682 0.901 0.779 0.984 2.229 855 6,885 2,364 170 0.753 0.834 0.789 0.976 2.598
1CT: Cheat type; 2Optimizer: Task-Specified Threshold Optimizer in Mvin; 3OEI: Operational Efficiency Index.
6.5 Hawk and Official Inspection Comparison
To answer RQ.2, we calculate the daily number of bans for the day and cumulative ban-sums by Hawk and the official inspections illustrated in Figure 6. The figure indicates that Hawk’s identification effectiveness (daily count in red bars) and efficiency (cumulative count in red lines) outrank those of the official manual inspections (blue bars and lines) on both cheat types. Additionally, Table 4 compares the average ban duration per match, revealing that Hawk’s seconds-scale processing time is exponentially more efficient than official inspections. The remediation rate indicates the percentage of cheaters not banned in the first labeling but caught in the second labeling. In Table 4, we list the remediation rate after the second labeling among Hawk’s false positives (which are obtained after the first labeling). It demonstrates that Hawk can capture the authentic cheaters among the normal players initially labeled by the official anti-cheats. This also suggests that while Hawk is robust to mislabeled data, using a purer dataset would further improve results.
@@ 6.5 Hawk and Official Inspection Comparison
6.6 Analysis on Different Mvin Settings
To address RQ.3, we conduct a comparative analysis of Hawk under different Mvin settings, as shown in Table 6. The Optimizer column indicates the priority choices of the Task-Specified Threshold Optimizer, such as F1-Score for highest F1-Score and $\text{Acc}_{\textit{rc}>0.75}$ for highest accuracy with recall above 75%. An inevitable trade-off exists between capturing more cheaters and minimizing false positives due to cheating sophistication and dataset noise (Section 6.2). As discussed in Section 5.4, GMs can dynamically adjust optimizers, allowing for quick retraining of Mvin, with new weight matrices and thresholds generated in about one minute. This adaptability aligns with the varying intensity of required anti-cheat measures, offering flexibility in balancing performance and efficiency. GMs can use different optimizers to meet specific anti-cheat needs. Interestingly, GMs do not always aim to catch all cheaters. They sometimes balance catching cheaters with permissiveness, as certain cheaters may bring short-term profits (e.g., in buyout games, where bans may be delayed until after refund periods). When complaints rise, GMs may then conduct a mass ban (a regularly organized operation to clean up the gaming environment) to restore community confidence. This also partially draws to the creation of Mvin.
6.7 Robustness to Cheat Evolution
@@
Although anti-cheat measures are often opaque to prevent circumvention, cheat developers treat them as black boxes, adapting and camouflaging their cheats on a timely basis to evade detection. Cheat evolution refers to the diminishing effectiveness of specific anti-cheat over time. Commercially used anti-cheat methods (e.g., signature or process detection, driver scanning, etc.) are susceptible to this evolution since they target specific cheats. However, Hawk starts the analysis from the behavior, thus cheaters should appear data-level distinctions mentioned in Section 2.2. Hence, no matter how the cheats adapt, as long as the cheaters still want to gain unfair advantages, it is hard to bypass the detection of Hawk. To address RQ.4, we test Hawk’s robustness to cheat evolution using the Aimbot dataset, divided into partitions (40 matches each). Classified through date, the dataset is divided into partitions (40 matches per partition). The reason for selecting the Aimbot dataset is that during its collection window, an official mass ban on aimbots took place (at the 10th partition), and the cheats are usually updated within a few days after such an incident. Figure 7 and Figure 8 demonstrate the ablated performances of subsets under different metrics of Hawk across different numbers of partitions for training. According to the average ban duration of the official manual inspections in Table 4, approximately three sequestration ban cycles have passed during the collection window, and at least two ban cycles after the mass ban occurred. The ablated results of Hawk on both sets have been fluctuating to maintain within a small range over partitions (time) and with the cheat evolution. Besides, in Figure 6, we observe the effectiveness of Hawk has not continuously decreased over time on both datasets (we separated and extended the collection windows in Wallhack dataset for further proof). Thus, both of the results indicate Hawk is robust to cheat evolution.
@@
Figure 7: Ablated robustness against cheat evolution over partitions on Aimbot’s validation set under different metrics.
Figure 8: Ablated robustness against cheat evolution over partitions on Aimbot’s test set under different metrics.
6.8 Detecting Uncaught Cheaters in the Wild
@@
To answer RQ.5, given the above-known presence of cheaters in the false positives in Table 4 after the first labeling, we conduct a use case analysis to explore if there are additional illegal activities that exist after the second labeling. Given the lengthy average duration of a match, which demands significant manpower, we randomly select 30 Hawk’s false positives from the Aimbot and Wallhack validation sets for manual review, respectively (experts’ skill level is listed in Table 8). With the assistance of three experts in FPS games, we identify 27 suspects, who are labeled as normal players after the second labeling, to be highly suspicious within these 60 false positives. Of these 27 suspects, 14 originate from the Aimbot dataset (4 are considered using aimbots and 10 are considered boosting, i.e., elevate the account rank by other higher-rank players playing the account), while 13 are from the Wallhack dataset (8 are considered using wallhacks and 5 are considered boosting). We reported these suspicions to the official personnel. Upon their review, they confirm all 12 of the aforementioned aimbots and wallhacks suspects are committed to cheats, subsequently issuing bans for these cheaters on the platform. This reaffirms the poor performance of current in-use anti-cheats and the datasets used for training and evaluation are not entirely pristine. The website mentioned in Section 1 demonstrates their illegal actions. In terms of boosting suspects, there are no existing solutions for the officials to evidently verify their illegal actions. Therefore, there are no bans issued on the suspicious boosting accounts.
@@
Table 7: Maximum overheads of Hawk’s subsystems per match. The numbers are retrieved by the Time [36] in Linux.

### v2

@@ VI-B Experimental Settings and Ground Truth
TABLE VIII: Hawk detection result with different Task-Specified Threshold Optimizer (TSTO) settings.
CT TSTO Validation Test
TP TN FP FN Accuracy Recall AUC-ROC NPV FPR TP TN FP FN Accuracy Recall AUC-ROC NPV FPR
Aimbot Balance1 389 7,390 307 475 0.909 0.450 0.705 0.940 0.040 457 7,444 719 427 0.873 0.517 0.714 0.946 0.088
F1-Score 448 7,305 392 416 0.906 0.519 0.734 0.946 0.051 548 6,928 1235 336 0.826 0.620 0.734 0.954 0.151
Accuracy 269 7,528 169 595 0.911 0.311 0.645 0.927 0.022 305 7,888 275 579 0.906 0.345 0.656 0.932 0.034
AUC-ROC 648 4,603 3,094 216 0.613 0.750 0.674 0.955 0.402 691 4,364 3,799 193 0.559 0.782 0.658 0.958 0.465
$\text{Acc}_{\textit{rc}>0.75}$2 652 5,321 2,376 212 0.698 0.755 0.723 0.962 0.309 672 5,365 2,798 212 0.667 0.760 0.709 0.962 0.343
$\text{Acc}_{\textit{rc}>0.8}$ 700 4,611 3,086 164 0.620 0.810 0.705 0.966 0.401 716 4,700 3,463 168 0.599 0.810 0.693 0.965 0.424
$\text{Acc}_{\textit{rc}>0.85}$ 736 3,790 3,907 128 0.529 0.852 0.672 0.967 0.508 761 3,875 4,288 123 0.512 0.861 0.668 0.969 0.525
$\text{Acc}_{\textit{rc}>0.9}$ 778 2,859 4,838 86 0.425 0.900 0.636 0.971 0.628 787 2,955 5,208 97 0.414 0.890 0.626 0.968 0.638
Wallhack Balance 592 8,256 489 375 0.911 0.612 0.778 0.957 0.056 616 8,344 905 409 0.872 0.601 0.752 0.953 0.098
F1-Score 576 8,313 432 391 0.915 0.596 0.773 0.955 0.049 651 8,040 1,209 374 0.846 0.635 0.752 0.956 0.131
Accuracy 374 8,582 163 593 0.922 0.387 0.684 0.935 0.019 167 9,158 91 858 0.908 0.163 0.577 0.914 0.010
$\text{Acc}_{\textit{rc}>0.7}$ 677 7,898 847 290 0.883 0.700 0.802 0.965 0.097 803 7,506 1,743 222 0.809 0.783 0.797 0.971 0.188
$\text{Acc}_{\textit{rc}>0.75}$ 727 7,589 1,156 240 0.856 0.752 0.810 0.969 0.132 859 6,964 2,285 166 0.761 0.838 0.795 0.977 0.247
$\text{Acc}_{\textit{rc}>0.8}$ 776 7,186 1,559 191 0.820 0.802 0.812 0.974 0.178 583 8,280 969 442 0.863 0.569 0.732 0.949 0.105
$\text{Acc}_{\textit{rc}>0.85}$ 822 6,712 2,033 145 0.776 0.850 0.809 0.979 0.232 871 6,810 2,439 154 0.748 0.850 0.793 0.978 0.264
$\text{Acc}_{\textit{rc}>0.9}$ 871 5,755 2,990 96 0.682 0.901 0.779 0.984 0.342 855 6,885 2,364 170 0.753 0.834 0.789 0.976 0.256
[…]
VI-C Overall Performance, Ablation and Baseline Comparison
To answer RQ.1, we conduct an ablation study, shown in Table IV. We set the TSTO to achieve the highest accuracy with at least 70% recall for detecting aimbot, and the highest AUC-ROC for detecting wallhack (same below unless otherwise stated), illustrating different anti-cheat scenarios (further evaluated in Section VI-F). Overall, Hawk excels in identifying authentic cheaters, evidenced by the highest recall and NPV across all evaluations. For aimbot detection, RevStats enhances accuracy by effectively filtering normal players in both validation and test sets. In wallhack detection, RevStats excels in the validation set, while ExSPC leads in the test set. Although some components like RevStats and ExSPC outperform Hawk in accuracy and FPR, their lower recall (up to 64%) limits their effectiveness, affirming Hawk as the optimal choice. Notably, the Wallhack dataset yields better overall performance than Aimbot. We believe this is due to the latter’s diverse subcategories of aimbots. We conduct the pair-wise ablation study to evaluate the performance among Hawk and different combinations of subsystems. In Table IV, we observe that RevPov and RevStats combination has significantly higher FPR on the Aimbot set and slightly lower recall on the Wallhack set. The other two combinations both suffer from lower recall. While evaluating the performance, we prioritize recall, accuracy and FPR on the test set over other metrics, because they can more directly reflect the ability to identify cheaters or the overall performance. Hawk’s framework is designed with both detection effectiveness and efficiency in mind. We select models with acceptable detection performance, and further select models with low training and prediction costs based on the overhead described in the relevant articles. Table VII and Table V show RevPov’s and RevStats’s promising model selection results among hundreds of tested baselines. RevPov outperforms other models with balanced, consistent accuracy and recall. For RevStats, we found that Random Forest outperformed other models. Yet, it sometimes failed to overcome the overfitting problem (e.g., wallhack validation set). Thus, we adopt multi-forest and multi-subsampling with Random Forest. As a result, RevStats outperforms individual classifiers in balancing recall and accuracy by capitalizing on the strengths of each classifier. We also prioritize metrics recall, accuracy and FPR on the test set over other metrics. For example, although Logistic Regression has a higher recall, its accuracy and FPR are significantly lower. The contribution of the recall’s increase can not make up for the decrease in accuracy and FPR.
The current framework has optimized overall performance compared with more innovative models. For example, in the preliminary test of RevPov’s encoder selection, we obtain lower detection performance with extended training and testing time while using the Transformer.
@@ VI-C Overall Performance, Ablation and Baseline Comparison
VI-D Hawk and Official Inspection Comparison
To answer RQ.2, we calculate the daily number of bans for the day and cumulative ban-sums by Hawk and the official inspections illustrated in Figure 7. The figure indicates that Hawk’s identification effectiveness (daily count in red bars) and efficiency (cumulative count in solid red lines) outrank those of the official manual inspections (hatched blue bars and dashed blue lines) on both cheat types. Additionally, Table VI compares the average ban duration per match, revealing that Hawk’s seconds-scale processing time is exponentially more efficient than official inspections. The remediation rate indicates the percentage of cheaters not banned in the first labeling but caught in the second labeling. Table VI lists the remediation rate after the second labeling among Hawk’s false positives (which are obtained after the first labeling). It demonstrates that Hawk can capture the authentic cheaters who escape from the official anti-cheat system. Moreover, this indicates that Hawk is robust against mislabeled data. We also compare the Hawk’s performance with the official in-use anti-cheat in the next section, as it is relevant to Mvin.
TABLE IX: Hawk comparative study results to prior works on the Aimbot test set. The optimal results are bolded.
Method Accuracy Recall NPV AUC-ROC FPR
AccA[11] 0.103 0.997 0.897 0.558 0.997
HitAcc[12] 0.675 0.545 0.931 0.658 0.311
Hawk 0.716 0.726 0.960 0.721 0.285
VI-E Comparative Analysis Among Hawk and Prior Works
To answer RQ.7, we compare Hawk with two other reproduced methods using Hawk’s Aimbot dataset. Due to the lack of available open-source solutions, we write our own code to reproduce both aimbot detection methods based on statistical features related to aiming or elimination used in prior works. Both methods are related to Hawk but use naive statistical features, data, and approach. Since prior work only targeted aimbots, we only used the Aimbot dataset for comparison. The results are listed in Table IX. Hawk outperforms prior works across all metrics but AccA’s recall, because AccA classifies most of the samples as cheaters. Below is a summary of each method:
  • AccA: Filters cheaters when the player’s field-of-view acceleration exceeds a threshold [11].
  • HitAcc: Filters cheaters based on the weapon accuracy of the player, using SVM [12].
VI-F Analysis on Different Mvin Settings
To address RQ.3, we conduct a comparative analysis of Hawk under different Mvin settings, as shown in Table VIII and footnote 4. In Table VIII, except Mvin, other subsystems are trained and validated with recall priority. By only adjusting the TSTO in Mvin, Hawk achieves rapid dynamic tuning ability, by using pre-trained Mvin or retraining Mvin to generate weight matrices and thresholds in about one minute. For example, the TSTO settings can be adjusted to achieve a lower FPR if necessary, avoiding the potential surge of excessive manual verification workload. Table VIII shows that leveraging different optimizer settings, e.g., Accuracy or F1-Score optimization modes, can obtain low FPR (1%-3% or 5%-15%, respectively). Alternatively, the Balance mode maintains a lower FPR than current industry standards (grey bar in footnote 4) while optimizing recall, resulting in slightly lower FPR and significantly higher recall compared to industry anti-cheat system (see blue bars in footnote 4, industry anti-cheats are introduced in Section VIII).
VI-G Robustness to Cheat Evolution
@@
Although anti-cheat measures are often opaque to prevent circumvention, cheat developers treat them as black boxes, adapting and camouflaging their cheats on a timely basis to evade detection. Cheat evolution refers to the diminishing effectiveness of specific anti-cheat over time. Commercially used anti-cheat methods are susceptible to this evolution (explained in Section VIII). However, Hawk starts the analysis from the behavior, thus cheaters must appear the data-level distinctions mentioned in Section II-B. Hence, no matter how the cheats adapt, as long as the cheaters still want to gain unfair advantages, it is hard to bypass the detection of Hawk. To address RQ.4, we test Hawk’s robustness to cheat evolution using the Aimbot dataset. Classified through date, the dataset is divided into partitions (40 matches per partition). The training and validation are performed on all data up to the current partition. We compared its performance on the same test set with different partitions’ trained models. The reason for selecting the Aimbot dataset is that during its collection window, an official mass ban on aimbots took place (at the 10th partition), and the cheats are usually updated within a few days after such an incident. 8(a) and 8(b) demonstrate the ablated performances of subsets under different metrics of Hawk across different numbers of partitions for training. According to the average ban duration of the official bans in Table VI, approximately three sequestration ban cycles have passed during the collection window, and at least two ban cycles after the mass ban occurred. This offers more than enough time for cheat to evolve. The ablated results of Hawk on both sets have been fluctuating to maintain within a small range over partitions (time) and with the cheat evolution. This indicates even without using the latter models of partitioned samples after the mass ban, Hawk ’s performance remains stable on the test set before and after the mass ban occurred. Besides, in Figure 7, we observe the effectiveness of Hawk has not continuously decreased over time on both datasets (we separated and extended the collection windows in Wallhack dataset for further proof). Thus, both experiments indicate Hawk is robust to cheat evolution.
(a) Validation set
(b) Test set
@@
Figure 8: Ablated robustness against cheat evolution under different metrics for validation (a) and test (b) sets.
VI-H Detecting Uncaught Cheaters in the Wild
@@
To answer RQ.5, given the above-known presence of cheaters in the false positives in Table VI after the first labeling, we conduct a use case analysis to explore if there are additional illegal activities that exist after the second labeling. Given the lengthy average duration of a match, which demands significant manpower, we select the top 30 Hawk’s false positives based on the decreasing order of the akpr from the Aimbot and Wallhack validation sets for manual review, respectively (experts’ skill level is listed in Supplemental Material Table 1). With the assistance of three experts in FPS games, we identify 27 suspects, who are labeled as normal players after the second labeling, to be highly suspicious within these 60 false positives. Of these 27 suspects, 14 originate from the Aimbot dataset (4 are considered using aimbots and 10 are considered boosting, i.e., elevate the account rank by other higher-rank players playing the account), while 13 are from the Wallhack dataset (8 are considered using wallhacks and 5 are considered boosting). We reported these suspicions to the official personnel. Upon their review, they confirm all 12 of the aforementioned aimbots and wallhacks suspects are committed to cheats, subsequently issuing bans for these cheaters. This reaffirms the poor performance of current in-use anti-cheats. The website mentioned in Section I demonstrates their illegal actions. In terms of boosting suspects, there are no existing solutions for the officials to evidently verify their illegal actions. Therefore, there are no bans issued on the suspicious boosting accounts. This experiment further proves the limited performance of current industrial anti-cheats, while demonstrating Hawk’s superior anti-cheat capabilities and high robustness to noise.
VI-I Overheads
Hawk is designed as an asynchronous, load-balancing system that analyzes post-game replay files during periods of low server activity. It operates independently of active game servers and can be deployed on a separate server. However, we still conduct overhead assessments for reference. To answer RQ.6, we utilize an Ubuntu server equipped with 128 cores AMD EPYC 7713P 64-Core Processor and 256 GB RAM, a real dedicated game server configuration [61]. Leveraging the pre-trained model in the artifact, Hawk requires only CPU resources to function efficiently. Table X shows the maximum overhead statistics per match, with each match analyzed concurrently. Additionally, we estimate the Hawk-to-server overheads in terms of a real deployment scenario. Our partner platform reports that at most 150,000 matches are played daily, with each server needing to handle up to 573 matches a day. According to our platform’s report and public server statistics [62], player activity is lowest between 1-5 AM local time, requiring analysis of up to 144 matches per hour. Based on the elapsed time in Table X, Hawk needs to process up to 12 matches concurrently to complete all analyses within the timeframe per server. We further conduct concurrent processing evaluation, it occupies at most 17.95% memory and 35.45% CPU usage for processing 12 matches concurrently. Therefore, each server has more than sufficient resources for deploying Hawk, even when hosting several matches simultaneously. It is vital for a design to maintain stable performance in high-concurrency environments [63]. To further validate the performance of Hawk under varying concurrency levels, we employed a game server equipped with a 24-core Intel® Xeon® Gold 5418Y CPUs and 503 GB RAM. During idle periods on this server, we concurrently request Hawk 10, 20, and 50 times, respectively. The results are listed in Table Table XI. The system maintains stable performance when handling different levels of concurrent requests. Furthermore, based on the aforementioned server load data, each server requires running Hawk concurrently 12 times per hour at most. Our tests reached a maximum concurrency of 50 requests, exceeding the data upper bound by over four times. But Hawk requires less than 7 minutes to complete all 50 requests. Therefore, we conclude that using Hawk under high concurrency conditions does not significantly affect the system overhead.
@@ VI-I Overheads
TABLE X: Maximum overheads of Hawk’s subsystems per match. The numbers are retrieved by the Time [64] in Linux.